please help me solve this equation -5/8h+25=15
Answer by stanbon(75887)

5/8h+25=15
----
Subtract 25 from both sides to get:
(5/8)h = -10
Multiply both sides by (8/5):
y = (8/5)(-10) = 8*-2 = -16
==================================
cheers,
Stan H.